Afreight train and an express train leave towns 390km apart, traveling toward one another. The freight train travels 30km per hr slower than the express train. They pass another 3 hr. later. What are their speeds!
Found 2 solutions by ichudov, noelsarrosa:
Answer by ichudov(499)

This is a typical travel problem. It asks us for their speeds, so, we need to find them! Denote speed of the express train as E, and speed of the freight train as F.
F=E-30
3*(F+E) = 390
OR
E-F=30
3E+3F = 390
it is a simple linear system, for which I have a solver that shows work.

Answer by noelsarrosa(1)

(x-30+x)3=390
x-30+x=130
2x-30=130
2x=160
x=80, x-30=50
Checking:
(80+50)3=390
130(3)=390
390=390